# Spatially resolved H emission in B14-65666: compact starbursts, ionizing efficiency and gas kinematics in an advanced merger at the Epoch of Reionization

### v1

@@ 4 Results and discussion
We detect and spatially resolve the rest-frame optical continuum and H emission in the system B14-65666 at â=â7.15 with MIRI imaging and spectroscopy. The new MIRI data are combined with ancillary NIRCam imaging and ALMA spectroscopy to obtain the physical properties of galaxies E and W, and of the entire system: star formation rate, instantaneous SFR, ionizing photon production efficiency, H equivalent width, dynamical mass, stellar mass surface density, and SFR surface density.
@@ 4.1 Spectral energy distribution and star formation history
We perform a SED fitting using CIGALE (Noll et al. 2009; Boquien et al. 2019) for galaxies E+W as well as for the two individual components. The SED covers the 1âm to 1.3âmm observed wavelength range, combining JWST photometric data (NIRCam and MIRIM) with ALMA continuum imaging (Hashimoto et al. 2019; Sugahara et al. 2021) along with our H and [OâIII]â88âm (Hashimoto et al. 2019) line fluxes. The redshift used in the analysis is the one derived spectroscopically from the H line.
We have made different assumptions to represent the star formation history of the system. The best results are determined from CIGALE fitting by minimizing the 2 value and ensuring the best agreement with observed photometry and line fluxes. These are obtained when considering two components: an instantaneous burst with ages ranging from 30 to 500 Myr, to model the mature stellar population, and a constant burst with ages lower than 10 Myr, to model the young stellar population. We use the Bruzual & Charlot (2003) stellar population models and the Chabrier (2003) Initial Mass Function (IMF). The nebular continuum and emission lines are modeled using electron densities from 100 to 1000 cm-3 and the ionization parameter (logU) ranging from -3.5 to -1.5. The metallicity is treated as a free parameter, ranging from 0.03âZâ to 0.7âZâ, as previous studies showed that it varies in the two galaxies E and W (Sugahara et al. 2025; Jones et al. 2024). We use the dust attenuation law from Calzetti et al. (2000), and the dust emission is modeled using the Draine et al. (2014) models.
The best-fit models for the galaxies E+W and for galaxies E and W are shown in Fig. 4. These models reveal that B14-65666 is dominated by a young stellar burst characterized with a total SFR of 127ââ66âMyr-1 and a mass for the young stellar component of (8ââ1)ââ108 Mâ. Galaxies E and W have SFRs of 84ââ49 and 50ââ12âMyr-1, respectively, and show the same young stellar component with a stellar mass of (4ââ1)ââ108âMâ. The SED fitting results indicate the presence of a mature stellar population in B14-65666 with a stellar mass of around 0.5ââ2ââ109âMâ and average ages about 200 Myr. SED fitting results are presented in Table 2. We conclude from the SED fitting that the B14-65666 system consists of two galaxies experiencing starbursts characterized by a similar age (see Sect. 4.9 for a more detailed discussion). Also, despite the compactness of galaxy E (see Sect. 4.7), the SED based on the NIRCam and MIRI imaging, and covering the rest-frame 0.16âââ0.78âm, does not correspond to that of the average Little Red Dots at redshifts â6âââ8 (PĂ©rez-GonzĂĄlez et al. 2024b; Kokorev et al. 2024; Rinaldi et al. 2024a).
@@ 4.2 Nebular and stellar attenuation
Dust attenuation plays a key role in shaping the observed properties of B14-65666, as revealed by MIRI, NIRCam, and ALMA observations. Combining the MRS H and the NIRSpec H fluxes444We implement aperture corrections (ACâ=â1/0.79) using the NIRSpec IFS PSF to match the fluxes presented in Jones et al. (2024) to the larger aperture of our MRS measurements., the H/H ratio gives observed values of 4.5â0.6 and 2.9ââ0.4 for galaxies E and W, respectively. Note that these values correspond to the narrow component of the line emission for both galaxies as the broad component present in the H line (Jones et al. 2024) is not detected in H with the present data (see Sect. 4.5). Assuming case B recombination, Teââ1.5âââK and neâ=ââcm-3, the theoretical ratio is 2.8. Using Cardelli et al. (1989), we get that the nebular visual extinction () is â=â1.5ââ0.4 mag for galaxy E, and it is â=â mag for galaxy W. Thus the nebular attenuation is higher for galaxy E, in agreement, within the uncertainties, with the values of â=â1.6ââ0.3 mag for galaxy E and â=â0.7â mag for W from Sugahara et al. (2025) (obtained with SED fitting and applying the conversion factor to nebular attenuation, fâ=â0.44; Calzetti 1997).
We also calculate the dust attenuation in the FUV, , with the IR-to-UV luminosity ratio (IRXâ=â/), as this is a robust tracer of . Taking the value of â=â0.7 for the total galaxy (Sugahara et al. 2025), we adopt expression (5) from Meurer et al. (1999) assuming the bolometric correction for the total light emitted by dust is 1, and for the total light emitted by stars, UV, is 1.68 (Meurer et al. 1999). We get a value of â=â1.50, compatible with that obtained with our SED fitting for galaxies E+W, â=â1.41â mag (Calzetti et al. 2000).
@@ 4.3 Star formation rate and burstiness
We derive the recent (i.e. less than 10 Myr) star formation rate directly from our H measurements, assuming solar (to facilitate comparison with other star formation tracers at other wavelengths) and sub-solar metallicity (0.29 Zâ, Jones et al. 2024 and see Table 2). We apply the corresponding SFR-to-H conversion factor for each metallicity, which is approximately 1.7 times lower in the subsolar case (Reddy et al. 2018). The final SFR values are corrected for attenuation using the H/H ratios (see previous section). We obtain SFRs of () Mââyr-1 and () Mââyr-1 for galaxies E and W and solar (subsolar) metallicities, respectively.
@@
Additional estimates of the SFR can be obtained directly from far-infrared emission lines, or using a combination of multi-wavelength tracers. The [CâII]â158â emission line is a tracer of SFR in low-redshift star-forming and composite SF+AGN galaxies (De Looze et al. 2014), confirmed for high-redshift galaxies (Schaerer et al. 2020; Carniani et al. 2020). Adopting the ancillary [CâII]â158âm luminosities derived from ALMA (Hashimoto et al. 2019), values similar to those derived from our H measurements are obtained, within the uncertainties (see Table 2). The large uncertainty in the value of the SFR derived from the [CâII]â158âm emission line is due to the nature of this emission which could include extended or diffuse regions not directly linked to active star formation. Also, to trace less massive stars, we use the SFR(UV), for solar metallicities and considering a constant SFH over 100 Myr (Calzetti 2013).
@@
Other SFR estimates involve the use of hybrid indicators such as the UV+TIR (Total Infrared) luminosity (Kennicutt & Evans 2012), or H+TIR (CatalĂĄn-Torrecilla et al. 2015; Calzetti 2013), that capture both obscured and unobscured star formation over longer periods of time (up to 100 Myr). Assuming a simple energy balance (Kennicutt et al. 2009), and considering the far-UV luminosity derived from our SED fitting (â=â(1.15ââ0.11)ââ), and the total infrared luminosity â=â(1.05ââ0.2)ââ (Hashimoto et al. 2019), we get SFR consistent with our extinction-corrected H estimates. The SFR derived from the H and TIR luminosities are also in good agreement with those obtained from UV+TIR (see Table 2). The SFRs derived in this analysis (70âââ140âMââyr-1) are lower than the SED-based SFR reported by Sugahara et al. (2025), SFRâ=â225ââMââyr-1, and the (narrow+broad) H-based SFR from Jones et al. (2024), SFRHÎČâ=â308ââ39âMââyr-1. This difference is attributed to variations in aperture sizes where the contribution of the extended emission (H) is significant, as well as other factors such as different estimates of the internal dust attenuation. Additionally, in the case of the H-based SFR, the presence of both narrow and broad components in the H line emission also contributes to the higher SFR value by a factor 2 (Jones et al. 2024).
@@
A useful diagnostic for assessing the recent star formation in high- galaxies is the ratio of the SFR(H)-to-SFR(UV), referred as the burstiness parameter and which probes star formation over the last 10âMyr relative to the last 100âMyr. If the ratios are above 1, it indicates that the galaxy is in a bursty phase (Atek et al. 2024). We get a value of the ratio SFR(H)-to-SFR(UV) for solar metallicities of 1.8ââ0.1 and 0.8ââ0.1 for galaxies E and W, respectively (we would expect similar values for subsolar metallicities). When compared to the range of burstiness ratios reported by Atek et al. (2024) (0.7âÂĄâzâÂĄâ1.5), values above or close to the unity are typical in low-mass galaxies (MâÂĄâMâ), and below unity in high-mass galaxies (MâÂżâMâ). Thus, our values suggest that galaxy E shows signs of bursty activity, forming stars at nearly twice the average rate of the past 100 Myr, it is in a mild bursty phase, while galaxy W is forming stars at a rate closer to this past 100 Myr, i.e. experiencing a more continuous burst. This aligns with the results from Navarro-Carrera et al. (2024), who analyzed H emitters across zââ2âââ6.5 using JADES data, observing a correlation between enhanced SFR(H)/SFR(UV) and EW0(H), reinforcing the scenario where galaxies with above-unity SFR(H)/SFR(UV) experience a bursty phase in the last âMyr.
4.4 Equivalent width of H and ionizing photon production efficiency
@@
To derive the rest-frame equivalent width of H (EW0(H)) for galaxies E+W and galaxies E and W, we make use of the F560W and H fluxes measured in our MIRIM and the MRS observations, respectively. First, we obtain rest-frame optical continuum emission by subtracting the contribution of the H line (Sect. 3.2) to the F560W flux (Sect. 3.3). The EW0(H) is then derived as the ratio of the H to the continuum flux under the H line. We obtain rest-frame equivalent widths 832ââ100âĂ and 536ââ78 Ă for galaxies E and W, respectively. The difference in the EW0(H) of the two galaxies is attributed to a different star formation history in the two galaxies with ionizing, young stars, and more mature, non-ionizing, stellar populations contributing in different proportions to the continuum emission. According to our SED analysis, galaxy E, with a larger EW0(H) value, has a slightly younger stellar population (6 Myr) relative to galaxy W (8 Myr).
@@
The ionizing photon production efficiency, ), is the ratio of the ionizing to the non-ionizing UV flux. Assuming recombination, ) is empirically obtained as the ratio of the number of ionizing photons () derived from the H line (after extinction correction) and the luminosity at 1500 Ă . Assuming zero escape fraction, a dust attenuation as derived in Sect. 4.2, and a temperature of 1.5â 104âK (Ălvarez-MĂĄrquez et al. 2024) the integrated H luminosity provides a total number of ionizing photons () equal to âphâs-1 and âphâs-1 for galaxies E and W, respectively. Different approaches to dust correction can yield significantly different values of introducing uncertainty in its determination and affecting the interpretation of its physical implications (Matthee et al. 2017). Thus, correcting the UV and H luminosities, we get ionizing photon production efficiencies ) equal to and Hz erg-1 for galaxies E and W, respectively. This difference in the between the two galaxies of the system can be interpreted as partly due to the age difference in the young stellar populations in the two galaxies. Models indicate that in addition to the stellar continuum, there is a strong nebular contribution due to the two-photon continuum at 1500. This contribution can be close to 50% of the total for very young populations (1 Myr), decreasing with age and becoming almost negligible for 10-Myr-old populations (Katz et al. 2024). Thus the impact of this nebular continuum would be to increase the 1500â luminosity and consequently decrease the corresponding by factors 0.15 to 0.04 for ages 1 to 8 Myr. In addition, residual differential extinction in the two galaxies could also play an effect due to uncertainties in the derivation of the extinction in the ultraviolet.
@@
We present the previous results, in the âââEW0(H) plane (see Fig. 5). Additional samples of galaxies at redshifts 4 to 9 are also presented for comparison. These include Lyman alpha emitters (LAE) at redshifts 3.8-6 (Maseda et al. 2020; Ning et al. 2023; Lam et al. 2019; Harikane et al. 2018), H and [OâIII] emitters at redshifts 6.1 to 8 (Sun et al. 2023; Rinaldi et al. 2023), and the mean of Lyman Break galaxies at redshift 8 (Stefanon et al. 2022). Individual galaxies with resolved integral field spectroscopy are also shown (RXCJ0600-2007 at â=â6.07, GimĂ©nez-Arteaga et al. 2024 and MACS1149-JD1 at â=â9.1, Ălvarez-MĂĄrquez et al. 2024). While galaxies E and W are forming stars at a rate much higher than the average galaxy at redshift 7, they have and EW0(H) in the range of high- galaxies. These values are consistent with the predicted values for young stellar populations. The Binary Population and Spectral Synthesis model (BPASS, Eldridge & Stanway 2020) demonstrates that the ) and the EW0(H) of a single-aged, instantaneous-starburst binary population increases as the age of the population decreases (see Fig. 5), with values compatible with those observed for young populations with ages from 1 to 20 Myr. The production of ionizing photons also depends on the presence of massive stars and therefore the initial mass function (Stanway et al. 2016). Values for the logarithm of the photon production efficiency above 25.5 Hzâerg-1 can only be obtained with massive bursts, âMâ due to the stochasticity of the formation of massive stars (Stanway & Eldridge 2023). The galaxies in the B14-65666 system and the high-z galaxies identified with JWST should have higher young stellar masses.

### v2

@@ 4 Results and discussion
We detected and spatially resolved the rest-frame optical continuum and H emission in the system B14-65666 at â=â7.15 with MIRI imaging and spectroscopy. The new MIRI data are combined with ancillary NIRCam imaging and ALMA spectroscopy to obtain the physical properties of galaxies E and W, and of the entire system: star formation rate, instantaneous SFR, ionizing photon-production efficiency, H equivalent width, dynamical mass, stellar-mass surface density, and SFR surface density.
@@ 4.1 Spectral energy distribution and star formation history
We performed an SED fitting using CIGALE (Noll et al. 2009; Boquien et al. 2019) for galaxies E+W as well as for the two individual components. The SED covers the 1âm to 1.3âmm observed wavelength range, combining JWST photometric data (NIRCam and MIRIM) with ALMA continuum imaging (Hashimoto et al. 2019; Sugahara et al. 2021), along with our H and [OâIII]â88âm (Hashimoto et al. 2019) line fluxes. The redshift used in the analysis is the one derived spectroscopically from the H line.
We made different assumptions to represent the star formation history of the system. The best results are determined from CIGALE fitting by minimizing the 2 value and ensuring the best agreement with observed photometry and line fluxes. These are obtained when considering two components: an instantaneous burst with ages ranging from 30 to 500 Myr to model the mature stellar population; and a constant burst with ages lower than 10 Myr to model the young stellar population. We used the Bruzual & Charlot (2003) stellar population models and the Chabrier (2003) initial mass function (IMF). The nebular continuum and emission lines were modeled using electron densities from 100 to 1000 cm-3 and the ionization parameter (logU) ranging from -3.5 to -1.5. The metallicity was treated as a free parameter, ranging from 0.03âZâ to 0.7âZâ, as previous studies showed that it varies in galaxies E and W (Sugahara et al. 2025; Jones et al. 2024). We used the dust attenuation law from Calzetti et al. (2000), and the dust emission was modeled using the Draine et al. (2014) models.
The best-fit models for galaxies E+W and for galaxies E and W separately are shown in Fig. 4. These models reveal that B14-65666 is dominated by a young stellar burst characterized with a total SFR of 127ââ66âMyr-1 and a mass for the young stellar component of (8ââ1)ââ108 Mâ. Galaxies E and W have SFRs of 84ââ49 and 50ââ12âMyr-1, respectively, and show the same young stellar component with a stellar mass of (4ââ1)ââ108âMâ. The SED fitting results indicate the presence of a mature stellar population in B14-65666 with a stellar mass of around 0.5ââ2ââ109âMâ and average ages about 200 Myr. SED fitting results are presented in Table 2. We conclude from the SED fitting that the B14-65666 system consists of two galaxies experiencing starbursts characterized by a similar age (see Sect. 4.9 for a more detailed discussion). Also, despite the compactness of galaxy E (see Sect. 4.7), the SED based on the NIRCam and MIRI imaging and covering the rest-frame 0.16âââ0.78âm does not correspond to that of the average little red dots at redshifts â6âââ8 (PĂ©rez-GonzĂĄlez et al. 2024b; Kokorev et al. 2024; Rinaldi et al. 2024a).
@@ 4.2 Nebular and stellar attenuation
Dust attenuation plays a key role in shaping the observed properties of B14-65666, as revealed by MIRI, NIRCam, and ALMA observations. Combining the MRS H and the NIRSpec H fluxes444We implemented aperture corrections (ACâ=â1/0.79) using the NIRSpec IFS PSF to match the fluxes presented in Jones et al. (2024) to the larger aperture of our MRS measurements., the H/H ratio gives observed values of 4.5â0.6 and 2.9ââ0.4 for galaxies E and W, respectively. We note that these values correspond to the narrow component of the line emission for both galaxies as the broad component present on the H line (Jones et al. 2024) is not detected in H with the present data (see Sect. 4.5). Assuming a case B recombination, Teââ1.5âââK and neâ=ââcm-3, the theoretical ratio is 2.8. Using Cardelli et al. (1989), we obtain that the nebular visual extinction () is â=â1.5ââ0.4 mag for galaxy E, and it is â=â mag for galaxy W. Thus the nebular attenuation is higher for galaxy E, in agreement, within the uncertainties, with the values of â=â1.6ââ0.3 mag for galaxy E and â=â0.7â mag for W from Sugahara et al. (2025) (obtained with SED fitting and applying the conversion factor to nebular attenuation, fâ=â0.44; Calzetti 1997).
We also calculated the dust attenuation in the FUV, , with the IR-to-UV luminosity ratio (IRXâ=â/), as this is a robust tracer of . Taking the value of â=â0.7 for the total galaxy (Sugahara et al. 2025), we adopted expression (5) from Meurer et al. (1999) assuming the bolometric correction for the total light emitted by dust is 1, and for the total light emitted by stars, UV, is 1.68 (Meurer et al. 1999). We obtain a value of â=â1.50, which is compatible with that obtained with our SED fitting for galaxies E+W, â=â1.41â mag (Calzetti et al. 2000).
@@ 4.3 Star formation rate and burstiness
We derived the recent (i.e., less than 10 Myr) star formation rate directly from our H measurements, assuming solar (to facilitate comparison with other star formation tracers at other wavelengths) and subsolar metallicity (0.29 Zâ, Jones et al. 2024 and see Table 2). We applied the corresponding SFR-to-H conversion factor for each metallicity, which is approximately 1.7 times lower in the subsolar case (Reddy et al. 2018). The final SFR values are corrected for attenuation using the H/H ratios (see previous section). We obtain SFRs of () Mââyr-1 and () Mââyr-1 for galaxies E and W and solar (subsolar) metallicities, respectively.
@@
Additional estimates of the SFR can be obtained directly from far-infrared emission lines or using a combination of multiwavelength tracers. The [CâII]â158â emission line is a tracer of SFR in low-redshift, star-forming, and composite SF+AGN galaxies (De Looze et al. 2014); it is confirmed for high-redshift galaxies (Schaerer et al. 2020; Carniani et al. 2020). Adopting the ancillary [CâII]â158âm luminosities derived from ALMA (Hashimoto et al. 2019), values similar to those derived from our H measurements are obtained, within the uncertainties (see Table 2). The large uncertainty in the value of the SFR derived from the [CâII]â158âm emission line is due to the nature of this emission, which could include extended or diffuse regions not directly linked to active star formation. Also, to trace less massive stars we used the SFR(UV) for solar metallicities and considering a constant SFH over 100 Myr (Calzetti 2013).
@@
Other SFR estimates involve the use of hybrid indicators such as the UV+TIR (total infrared) luminosity (Kennicutt & Evans 2012), or H+TIR (CatalĂĄn-Torrecilla et al. 2015; Calzetti 2013), that capture both obscured and unobscured star formation over longer periods of time (up to 100 Myr). Assuming a simple energy balance (Kennicutt et al. 2009), and considering the far-UV luminosity derived from our SED fitting (â=â(1.15ââ0.11)ââ), and the TIR luminosity â=â(1.05ââ0.2)ââ (Hashimoto et al. 2019), we obtain an SFR consistent with our extinction-corrected H estimates. The SFR derived from the H and TIR luminosities are also in good agreement with those obtained from UV+TIR (see Table 2). The SFRs derived in this analysis (70âââ140âMââyr-1) are lower than the SED-based SFR reported by Sugahara et al. (2025), SFRâ=â225ââMââyr-1, and the (narrow+broad) H-based SFR from Jones et al. (2024), SFRHÎČâ=â308ââ39âMââyr-1. This difference is attributed to variations in aperture sizes where the contribution of the extended emission (H) is significant, as well as other factors such as different estimates of the internal dust attenuation. Additionally, in the case of the H-based SFR, the presence of both narrow and broad components in the H line emission also contributes to the higher SFR value by a factor of 2 (Jones et al. 2024).
@@
A useful diagnostic for assessing the recent star formation in high- galaxies is the ratio of the SFR(H)-to-SFR(UV), which is referred to as the burstiness parameter and probes star formation over the last 10âMyr relatively to the last 100âMyr. If the ratios are above 1, it indicates that the galaxy is in a bursty phase (Atek et al. 2024). We obtain a value of the ratio SFR(H)-to-SFR(UV) for solar metallicities of 1.8ââ0.1 and 0.8ââ0.1 for galaxies E and W, respectively (we would expect similar values for subsolar metallicities). When compared to the range of burstiness ratios reported by Atek et al. (2024) (0.7âÂĄâzâÂĄâ1.5), values above or close to unity are typical in low-mass galaxies (MâÂĄâMâ) and below unity in high-mass galaxies (MâÂżâMâ). Thus, our values suggest that galaxy E shows signs of bursty activity, forming stars at nearly twice the average rate of the past 100 Myr; it is in a mild bursty phase, while galaxy W forms stars at a rate closer to that if this past 100 Myr; i.e., experiencing a more continuous burst. This aligns with the results from Navarro-Carrera et al. (2024), whose authors analyzed H emitters across zââ2âââ6.5 using JADES data, observing a correlation between enhanced SFR(H)/SFR(UV) and EW0(H), reinforcing the scenario where galaxies with above-unity SFR(H)/SFR(UV) experience a bursty phase in the last âMyr.
4.4 Equivalent width of H and ionizing photon-production efficiency
@@
To derive the rest-frame equivalent width of H (EW0(H)) for galaxies E+W and galaxies E and W, we make use of the F560W and H fluxes measured in our MIRIM and the MRS observations, respectively. First, we obtain rest-frame optical continuum emission by subtracting the contribution of the H line (Sect. 3.2) to the F560W flux (Sect. 3.3). The EW0(H) is then derived as the ratio of the H to the continuum flux under the H line. We obtain rest-frame equivalent widths of 832ââ100âĂ and 536ââ78 Ă for galaxies E and W, respectively. The difference in the EW0(H) of the two galaxies is attributed to a different star formation history in the two galaxies with ionizing young stars and more mature, non-ionizing stellar populations contributing in different proportions to the continuum emission. According to our SED analysis, galaxy E, with a larger EW0(H) value, has a slightly younger stellar population (6 Myr) relative to galaxy W (8 Myr).
@@
The ionizing photon production efficiency, ), is the ratio of the ionizing to the non-ionizing UV flux. Assuming recombination, ) is empirically obtained as the ratio of the number of ionizing photons () derived from the H line (after extinction correction) and the luminosity at 1500 Ă . Assuming zero escape fraction, a dust attenuation as derived in Sect. 4.2, and a temperature of 1.5â 104âK (Ălvarez-MĂĄrquez et al. 2024) the integrated H luminosity provides a total number of ionizing photons () equal to âphâs-1 and âphâs-1 for galaxies E and W, respectively. Different approaches to dust correction can yield significantly different values of introducing uncertainty in its determination and affecting the interpretation of its physical implications (Matthee et al. 2017). Thus, correcting the UV and H luminosities, we obtain ionizing photon-production efficiencies ) equal to and Hz erg-1 for galaxies E and W, respectively. This difference in the between the two galaxies of the system can be interpreted as partly due to the age difference in the young stellar populations in the two galaxies. Models indicate that in addition to the stellar continuum, there is a strong nebular contribution due to the two-photon continuum at 1500. This contribution can be close to 50% of the total for very young populations (1 Myr), decreasing with age and becoming almost negligible for 10-Myr-old populations (Katz et al. 2024). Thus, the impact of this nebular continuum would be to increase the 1500â luminosity and consequently decrease the corresponding by factors of 0.15 to 0.04 for ages of 1 to 8 Myr. In addition, residual differential extinction in the two galaxies could also play a role due to uncertainties in the derivation of the extinction in the UV.
@@
We present the previous results in the âââEW0(H) plane (see Fig. 5). Additional samples of galaxies at redshifts 4 to 9 are also presented for comparison. These include Lyman-alpha emitters (LAE) at redshifts 3.8-6 (Maseda et al. 2020; Ning et al. 2023; Lam et al. 2019; Harikane et al. 2018), H and [OâIII] emitters at redshifts 6.1 to 8 (Sun et al. 2023; Rinaldi et al. 2023), and the mean of Lyman-break galaxies at redshift 8 (Stefanon et al. 2022). Individual galaxies with resolved integral-field spectroscopy are also shown (RXCJ0600-2007 at â=â6.07, GimĂ©nez-Arteaga et al. 2024 and MACS1149-JD1 at â=â9.1, Ălvarez-MĂĄrquez et al. 2024). While galaxies E and W forms stars at a rate much higher than the average galaxy at redshift 7, they have and EW0(H) in the range of high- galaxies. These values are consistent with the predicted values for young stellar populations. The Binary Population and Spectral Synthesis (BPASS; Eldridge & Stanway 2020) model demonstrates that the ) and the EW0(H) of a single-aged, instantaneous-starburst binary population increases as the age of the population decreases (see Fig. 5), with values compatible with those observed for young populations with ages from 1 to 20 Myr. The production of ionizing photons also depends on the presence of massive stars and therefore the IMF (Stanway et al. 2016). Values for the logarithm of the photon production efficiency above 25.5 Hzâerg-1 can only be obtained with massive bursts, âMâ due to the stochasticity of the formation of massive stars (Stanway & Eldridge 2023). The galaxies in the B14-65666 system and the high-z galaxies identified with the JWST should have higher young stellar masses.